OpenStudy (anonymous):

What is the square root of -88?

OpenStudy (chaise):

88i

OpenStudy (anonymous):

It will have a imaginary root

OpenStudy (anonymous):

\[i \times 2 \sqrt{22}\]

OpenStudy (anonymous):

I had 88i, but I'm not sure if that's what my teacher is looking for. The square root of -169 simplified to 13i, so maybe the -88 simplifies more?

OpenStudy (chaise):

yes, it simplifies to \[\sqrt{4}\times \sqrt{22}\] The square root of 4 is 2. so the answer is 2(sqrt(22))i Do you understand this?

OpenStudy (anonymous):

How did you get 2 and 22?

OpenStudy (anonymous):

I mean 4 and 22.

OpenStudy (anonymous):

\[88 = 2 \times 2\times 2 \times 11\]

OpenStudy (chaise):

You can split up square roots like this: sqrt(a) x sqrt(b) = sqrt(ab) Does this make sense as to how I got 4 and 22?